[4/3*9/2]+3/2= Simplify. I started breaking it down and I came up with 4/3*9/2=36/6=6 6+3/2= I don't know the next step. When I put this problem in my calculator it gives me 15/2 and I don't know how it got to that answer. What am I forgetting?
Answer by anjulasahay(30)

[4/3*9/2]+3/2= Simplify. I started breaking it down and I came up with 4/3*9/2=36/6=6 6+3/2= I don't know the next step. When I put this problem in my calculator it gives me 15/2 and I don't know how it got to that answer. What am I forgetting?
ans:
6+3/2=(2*6+3)/2 = 15/2